# Bootstrapping Classical Shadows for Neural Quantum State Tomography

### v1

@@ I Introduction
Recently, Ref. [2] proposed an approach for training a neural quantum state using classical shadows to combine the advantages of variational ansatze and classical shadow measurements. The authors introduce the infidelity between the classical shadow state (introduced in [1]) and the ansatz as their loss function, enabling the training of the ansatz on Clifford measurements. They suggest that, unlike the traditionally used cross-entropy loss function, infidelity offers a better optimization landscape due to its metric properties and bounded range. However, in the case of an unphysical state like the classical shadow state, the infidelity is not constrained within the physically valid $[0,1]$ range. Additionally, infidelity cannot guarantee bounded errors in approximating many quantities of interest from the trained model. Finally, the loss does not generalize to ansatze for mixed states.
Moreover, Ref. [2] faces many practical challenges in its training protocol. The trainability of the authors’ ansatz relies on an ad hoc initialization step, without which the infidelity-based optimization stalls. To overcome this barren plateau, they pretrain their model using the cross-entropy loss function on the classical dataset with measurements only in the computational basis. This pretraining requires an additional set of measurements beyond the one used in the classical shadow technique. Furthermore, its success depends on the informational content of the underlying quantum state in the computational basis, rendering it non-universal.
@@
Results on pure states. To showcase the strengths of our ideas, we focus on the Greenberger–Horne–Zeilinger (GHZ) state that is typically used as a challenging benchmark for quantum state tomography due to its multi-partite entanglement. In Section III.1, we employ Clifford measurements of pure GHZ states to conduct a comparative study. Our new loss function and importance sampling strategy show superior performance in learning six- and eight-qubit GHZ states compared to existing methods [2, 3].
@@ II.3.1 Shadow-Based Loss Functions
Here, $p_{\lambda}(\phi_{i})=\lvert\innerproduct{\psi_{\lambda}}{\phi_{i}}\rvert^{2}$ is the probability of measuring the snapshot $\phi_{i}=U_{i}|s_{i}\rangle$.
The same loss function can also be modified for the case of Pauli ensembles as follows. For each Pauli $U_{i}=U_{i,1}\otimes\cdots\otimes U_{i,n}$ and any bitstring $b\in\{0,1\}^{n}$, let $|b|=\sum_{i=1}^{n}b_{i}$ be the number of ones in $b$. Let $U_{i,b}=U_{i,1}^{b_{1}}\otimes\cdots\otimes U_{i,n}^{b_{n}}$ be the operator that has an identity factor for each zero index in $b$. In this case,
@@
$\mathcal{L}_{\rm{inf}}^{\rm{P}}(\lambda)=-\sum_{i=1}^{N}\sum_{b\in\mathcal{B}_{n}}\sum_{c\in\mathcal{B}_{n-|b|}}3^{|b|}p_{\lambda}(\phi_{i,b,c}).$ (4)
@@
Let $\mathcal{D}\subseteq\rm{Stab}$ be an independent identically distributed family of samples (allowing repetitions) drawn from $p^{\mathrm{data}}$, and $\widetilde{\mathcal{D}}$ be the set of elements in $\mathcal{D}$ (removing repetitions). The cross-entropy loss function can be approximated by the log-likelihood of $\mathcal{D}$,
@@
$p^{\mathrm{emp}}(\phi)=\frac{1}{|\mathcal{D}|}\sum_{\phi^{\prime}\in\mathcal{D}}\mathbb{I}_{\phi^{\prime}=\phi}\,,$
@@ II.3.2 Monte Carlo Estimation and Gradients
Central to the computational tractability of the considered loss functions Eqs. 3, 5 and 7 is the evaluation of the overlap between a stabilizer state $\phi$ and the neural state $\psi_{\lambda}$:
$\innerproduct{\psi_{\lambda}}{\phi}=\sum_{s\in\mathcal{B}}\psi_{\lambda}^{*}(s)\phi(s).$
@@ III Results
III.1 Benchmarking Loss Functions on Pure GHZ States
@@
Figure 1: Benchmarking results of the performance of four methods in reconstructing both six- and eight-qubit GHZ states. We conduct our experiments using 1000 classical shadows based on Pauli or Clifford measurements (as specified in the titles above the plots), averaging the results over five training trials on the same dataset trained over 50 epochs. Each estimate of the overlap between a neural quantum state and a stabilizer (see Section II.3.2) is based on 500 newly generated samples. The final states for each loss function are represented by markers of a corresponding colour. Circular and square markers represent the samples from the last epoch for the stabilizer- and neural state-based loss functions, respectively. The upper row displays the loss functions while the lower row shows the associated infidelity with respect to the true state as a measure of the generalization power of the neural quantum state.
We illustrate the trainability and the generalization power of these training methods in Fig. 1 via the optimization progress curves of loss functions evaluated on the training set (upper row) and infidelity with respect to the true states (lower row), over 50 epochs. For Pauli measurements, plotted in Fig. 1(a) and (b), the superiority of cross-entropy-based training over the infidelity-based approach is apparent. The latter struggles to achieve significant progress in approximating the underlying quantum state. Moreover, the benefits of the cross-entropy-based loss function become even more pronounced when considering computational costs. Both loss functions rely on computing the overlap between the neural quantum state and stabilizers (see Section II.3.2). For the infidelity loss function, each Pauli measurement leads to a proliferation of stabilizers that grows super-exponentially with the number of qubits (see Eq. 4). In contrast, this prohibitive scaling is entirely absent for both cross-entropy-based loss functions, as each Pauli measurement contributes a single stabilizer. In a previous work [2], we attempted to address the prohibitive scaling of the infidelity-based approach through Monte Carlo sampling of stabilizers from a classical shadow state. However, this strategy led to unstable training. Additionally, we note that stabilizer-based training improves in-training performance across all loss functions (the results for neural-state-based cross-entropy loss functions have been omitted for visual clarity). However, these improvements during training translate to enhanced generalization only for cross-entropy loss functions. In fact, for the cross-entropy loss function, stabilizer-based training proves to be crucial.
@@
III.2 Applications on Mixed GHZ States
@@ III.2.1 Pauli Observables
III.2.2 Purity and Trace Distance
@@ Appendix A The Transformer Neural Quantum State and Details of Numerical Experiments
Table 1: Numerical experiment settings for Figs. 1, 2 and 3.
@@ Acknowledgements
We thank our editor, Marko Bucyk, for his careful review and editing of the manuscript. W. K. acknowledges the support of Mitacs and a scholarship through the Perimeter Scholars International program. P. R. acknowledges the financial support of Mike and Ophelia Lazaridis, Innovation, Science and Economic Development Canada (ISED), and the Perimeter Institute for Theoretical Physics. Research at the Perimeter Institute is supported in part by the Government of Canada through ISED and by the Province of Ontario through the Ministry of Colleges and Universities.

### v2

@@ I Introduction
Recently, Ref. [2] proposed an approach for training a neural quantum state using classical shadows to combine the advantages of variational ansatze and classical shadow measurements. The authors introduce the infidelity between the classical shadow state (introduced in Ref. [1]) and the ansatz as their loss function, enabling the training of the ansatz on Clifford measurements. They suggest that, unlike the traditionally used cross-entropy loss function, infidelity offers a better optimization landscape due to its metric properties and bounded range. However, in the case of an unphysical state like the classical shadow state, the infidelity is not constrained within the physically valid $[0,1]$ range. Additionally, infidelity cannot guarantee bounded errors in approximating many quantities of interest from the trained model. Finally, the loss does not generalize to ansatze for mixed states.
Moreover, Ref. [2] faces important practical challenges in its training protocol. The trainability of the authors’ ansatz relies on an ad hoc initialization step. That is, they pretrain their model using the cross-entropy loss function on the classical dataset with measurements only in the computational basis, which was specifically chosen so for the GHZ state. This pretraining requires an additional set of measurements beyond the one used in the classical shadow technique. Furthermore, its success depends on the informational content of the underlying quantum state in the computational basis, rendering it non-universal.
@@
Results on pure states. To showcase the strengths of our ideas, we focus on the Greenberger–Horne–Zeilinger (GHZ) state that is typically used as a challenging benchmark for quantum state tomography due to its multi-partite entanglement. In Section III.2, we employ Clifford measurements of pure GHZ states to conduct a comparative study. Our new loss function and importance sampling strategy show superior performance in learning six- and eight-qubit GHZ states compared to existing methods [2, 3].
@@ II.3.1 Shadow-Based Loss Functions
Here, “C” stands for Clifford ensembles, “inf” refers to the infidelity, and $p_{\lambda}(\phi_{i})=\lvert\innerproduct{\psi_{\lambda}}{\phi_{i}}\rvert^{2}$ is the probability of measuring the snapshot $\phi_{i}=U_{i}|s_{i}\rangle$.
The same loss function can also be modified for the case of Pauli ensembles as follows. For each Pauli $U_{i}=U_{i,1}\otimes\cdots\otimes U_{i,n}$ and any bitstring $b\in B_{n}\equiv\{0,1\}^{n}$, let $|b|=\sum_{i=1}^{n}b_{i}$ be the number of ones in $b$. Let $U_{i,b}=U_{i,1}^{b_{1}}\otimes\cdots\otimes U_{i,n}^{b_{n}}$ be the operator that has an identity factor for each zero index in $b$. In this case,
@@
$\mathcal{L}_{\rm{inf}}^{\rm{P}}(\lambda)=-\sum_{i=1}^{N}\sum_{b\in\mathcal{B}_{n}}\sum_{c\in\mathcal{B}_{n-|b|}}3^{|b|}p_{\lambda}(\phi_{i,b,c}),$ (4)
where “P” stands for Pauli ensembles.
@@
Let $\mathcal{D}\subseteq\rm{Stab}$ be an independent identically distributed family of samples (allowing repetitions) drawn from $p^{\mathrm{data}}$, and $\widetilde{\mathcal{D}}$ be the set of elements in the dataset $\mathcal{D}$ (removing repetitions). The cross-entropy loss function can be approximated by the log-likelihood of $\mathcal{D}$,
@@
$p^{\mathrm{emp}}(\phi)=\frac{1}{|\mathcal{D}|}\sum_{\phi^{\prime}\in\mathcal{D}}\mathbb{I}_{\phi^{\prime}=\phi}\,,$ (6)
@@ II.3.2 Monte Carlo Estimation and Gradients
Central to the computational tractability of the considered loss functions Eqs. 3, 5 and 8 is the evaluation of the overlap between a stabilizer state $\phi$ and the neural state $\psi_{\lambda}$:
$\innerproduct{\psi_{\lambda}}{\phi}=\sum_{s\in\mathcal{B}}\psi_{\lambda}^{*}(s)\phi(s).$ (9)
@@ III Results
III.1 Scaling and Variance Study
Prior to studying the training of neural quantum states with our new methods, we explore the benefits of using shadow weights and stabilizer-based importance sampling. To this end, we focus on the key training metrics: the informational content and the statistical properties of the gradient estimator.
To facilitate our analysis, we consider an empirical proxy for the data distribution $p^{\rm{data}}_{\rm{emp}}(\phi)\propto p^{\rm{data}}(\phi)\mathbb{I}_{\phi\in\widetilde{\mathcal{D}}}$, where $\mathbb{I}_{\phi\in\widetilde{\mathcal{D}}}$ is the indicator function that excludes unobserved stabilizers. This distribution represents the most-accurate reconstruction of the probability measure $p^{\rm{data}}$ given a dataset $\mathcal{D}$, assuming no prior knowledge about $p^{\rm{data}}$. To assess the informational content in the shadow weights, we consider the KL divergence ${\rm{KL}}\left(p^{\rm{data}}_{\rm{emp}}||p^{\rm{target}}\right)$ between this typically unknown distribution and $p^{\rm{target}}$, which denotes either of the two empirically accessible distributions used for training: $p^{\rm{emp}}$ based on Eq. 6 or $p^{\rm{sh}}$ specified by Eq. 7. Specifically, we study the dependence of ${\rm{KL}}\left(p^{\rm{data}}_{\rm{emp}}||p^{\rm{target}}\right)$ on the size of the dataset relative to the size of the system by scaling the number of qubits given a fixed dataset, as shown in Fig. 1(a), and scaling the number of measurements in the dataset for a fixed number of qubits, as depicted in Fig. 1(b).
Figure 1: Analysis of the informational content within shadow weights through the KL divergence ${\rm{KL}}\left(p^{\rm{data}}_{\rm{emp}}||p^{\rm{target}}\right)$. The KL divergence serves as a comparison of the proxy for the typically unknown data distribution $p^{\rm{data}}_{\rm{emp}}$ (see the main text for the definition) with the distributions that can be inferred from the shadow measurements $p^{\rm{emp}}$ based on Eq. 6 and $p^{\rm{sh}}$ specified by Eq. 7. The KL divergence is examined as a function of (a) the system size, i.e., the number of qubits, for a dataset of 1000 shadows and (b) the number of shadows for a six-qubit system. The measurement type is specified by the marker type: circles and squares for Pauli (indicated by the subscript “P”) and Clifford (“C”) measurements, respectively. The target distribution type is specified by the colours and line styles: blue dashed and orange solid curves for the empirical distribution and shadow weights, respectively. Our results are averaged over 32 instances of pure states produced using random Clifford circuits, with error bars indicating the standard deviation.
The advantage of shadow weights in terms of estimating $p^{\rm{data}}_{\rm{emp}}$ is evident for smaller systems, as shown in Fig. 1 (a), though it gradually diminishes as the systems grow larger. In contrast, Fig. 1(b) indicates that this benefit can be reclaimed by increasing the number of measurements performed. A crucial factor in interpreting the plots in Fig. 1 is considering the difference in the configuration space dimensions of $p^{\rm{data}}_{\rm{emp}}$ associated with Pauli and Clifford measurements. Notably, Pauli measurements can reconstruct $4^{n}$ distinct stabilizers, whereas Clifford measurements can reconstruct $2^{0.5+\mathcal{O}(1)n^{2}}$ distinct stabilizers [aaronson2004improved]. For $n=6$ qubits, the dataset sizes considered in Fig. 1(b) must be compared to the dimension of the stabilizer space. For the Pauli distirbution, the dataset ranges from 5% to 500% of the total set size of distinct stabilizers. However, for the Clifford distribution, the maximum number of measurements is less than $0.001\%$ of the corresponding stabilizer set size. This significant difference explains the slight scaling advantage of the $p^{\rm{emp}}$ Pauli measurements over the $p^{\rm{emp}}$ Clifford measurements, as the former have a higher likelihood of obtaining repeated stabilizers, while the latter degenerate to a uniform distribution. For shadow weights, we observe a scaling advantage over empirical distributions in both regimes. However, unlike in the case of the empirical distributions, the Clifford distribution shows a slight improvement over the Pauli distribution. Such an advantage of Clifford shadow weights over Pauli ones, despite being in a regime with less data, is consistent with the theoretical expectation of an exponential benefit in predicting high-weight, low-rank observables (i.e., other stabilizers) [1]. Therefore, given a sufficiently large measurement dataset, training with shadow weights-based cross-entropy is expected to produce neural quantum states that better capture the underlying quantum state.
We now evaluate the advantages of the proposed stabilizer-based method of Section II.3.2, over the neural state-based method of Eq. 10, for Monte Carlo estimation of the overlap, which is used to compute the training gradient. For benchmarking purposes, we analyze the errors in the gradient estimates relative to the true gradient, obtained through an exhaustive evaluation of the inner product Eq. 9. Specifically, we focus on the error in the gradient direction and distinguish between errors calculated over a mini-batch and a full batch of measurements. Figure 2 illustrates the evolution of these errors as the neural quantum state is trained using the true gradient. Figure 2(a) shows that both estimators struggle most with capturing the correct gradient direction during the early stages of training, where the neural quantum state is likely the most different from the target state. However, after several epochs, as the neural state becomes more similar to the target, the stabilizer-based estimation significantly improves, while the neural state estimation remains relatively high and inconsistent. Throughout the entire training process, the stabilizer-based estimator consistently provides estimates with much lower variance, which is essential for training stability. Additionally, our proposed estimator is consistently more accurate than the neural-based estimator. Being an unbiased estimator, the stochastic error due to mini-batching is expected to be cancelled out when averaged over the mini-batches and thus can be lowered using smaller learning rates, albeit at the cost of slower training dynamics. Figure 2(b) shows whether this expected behaviour holds for the given estimators. There is a clear difference between them: the stabilizer-based estimator functions as an unbiased estimator, reducing gradient errors by averaging over mini-batches. In contrast, the neural state-based estimator exhibits a systematic bias throughout the entire training duration, resulting in an extreme accumulation of errors from the true gradient. To eliminate this finite-sample bias, a significant increase in the samples generated by the neural state would be necessary. Therefore, given a fixed computational resource budget, our benchmarking results demonstrate that the stabilizer-based gradient estimation has a significant advantage over the neural state-based gradient estimation, highlighting its potential for optimization.
Figure 2: Discrepancy in angle for the Monte Carlo-estimated gradients compared to the exact gradient throughout the training process. The dataset consists of 1000 Clifford measurements performed on a three-qubit GHZ state. Estimation using neural state-based sampling (NS) and stabilizer-based sampling (SS) are depicted in blue and orange, respectively. Each estimation uses 500 Monte Carlo samples. The gradient errors are evaluated over (a) mini-batches and (b) a full batch. In the case of the mini-batch gradient error, the error bars correspond to the standard deviation in the error over all mini-batches.
III.2 Training Performance on Pure GHZ States
@@
Figure 3: Benchmarking results of the performance of four methods in reconstructing both six- and eight-qubit GHZ states. We conduct our experiments using 1000 classical shadows based on Pauli or Clifford measurements (as specified in the titles above the plots), averaging the results over five training trials on the same dataset trained over 50 epochs. Each estimate of the overlap between a neural quantum state and a stabilizer (see Eq. 9) is based on 500 newly generated samples. The final states for each loss function are represented by markers. Triangle and square markers represent the samples from the last epoch for the stabilizer- and neural state-based infidelity loss functions, respectively. Circular and “x” markers represent the samples from the last epoch for the empirical and shadow based cross-entropy loss functions, respectively. The upper row displays the loss functions while the lower row shows the associated infidelity with respect to the true state as a measure of the generalization power of the neural quantum state.
We illustrate the trainability and the generalization power of these training methods in Fig. 3 via the optimization progress curves of loss functions evaluated on the training set (upper row) and infidelity with respect to the true states (lower row), over 50 epochs. For Pauli measurements, plotted in Fig. 3(a) and (b), the superiority of cross-entropy-based training over the infidelity-based approach is apparent. The latter struggles to achieve significant progress in approximating the underlying quantum state. Moreover, the benefits of the cross-entropy-based loss function become even more pronounced when considering computational costs. Both loss functions rely on computing the overlap between the neural quantum state and stabilizers (see Eq. 9). For the infidelity loss function, each Pauli measurement leads to a proliferation of stabilizers that grows super-exponentially with the number of qubits (see Eq. 4). In contrast, this prohibitive scaling is entirely absent for both cross-entropy-based loss functions, as each Pauli measurement contributes a single stabilizer. In a previous work [2], we attempted to address the prohibitive scaling of the infidelity-based approach through Monte Carlo sampling of stabilizers from a classical shadow state. However, this strategy led to unstable training. Additionally, we note that stabilizer-based training improves in-training performance across all loss functions (the results for neural-state-based cross-entropy loss functions have been omitted for visual clarity). Moreover, these improvements during training translate to enhanced generalization only for cross-entropy loss functions. In fact, for the cross-entropy loss function, stabilizer-based training proves to be crucial.
@@
III.3 Applications on Mixed GHZ States
@@ III.3.1 Pauli Observables
III.3.2 Purity and Trace Distance
@@ Appendix A The Transformer Neural Quantum State and Details of Numerical Experiments
Table 1: Numerical experiment settings for Figs. 3, 4 and 5.
@@ Acknowledgements
We thank our editor, Marko Bucyk, for his careful review and editing of the manuscript. The authors acknowledge Bill Coish and Christine Muschik for useful discussions. W. K. acknowledges the support of Mitacs and a scholarship through the Perimeter Scholars International program. P. R. acknowledges the financial support of Mike and Ophelia Lazaridis, Innovation, Science and Economic Development Canada (ISED), and the Perimeter Institute for Theoretical Physics. Research at the Perimeter Institute is supported in part by the Government of Canada through ISED and by the Province of Ontario through the Ministry of Colleges and Universities.